The difference between heats of reaction at constant pressure and constant volume for the reaction


2C6H6 (I) + 1502(g) "12CO2 (g) + 6H2O (I)


at 25?C in kJ mol-1 is




1)


 -7.43 





2)


 +3.72





3)


-3.72    





4)


+7.43




Grade:12

2 Answers

Askiitians Expert Sahil Arora - IITD
13 years ago

Dear Student,

Heat of reaction at constant pressure is ΔH and at constant volume is ΔU.

ΔH-ΔU=ΔngRT

Δng=12-15=-3

ΔngRT=(-3)*(8.314*10-3KJ/mol/K)*(298 K)=-7.43

one more thing,someone may think of dividing the reaction by 2..but that will be wrong.

On multiplying by a factor,ΔH and ΔU also get multiplied by the factor..

All the best Pratik !!!




Regards,

Askiitians Experts

Sahil Arora
Saurabh Kumawat
5 years ago
2C6H6(l)+15O2(g) -> 12CO2(g)+6H2O(l) at 25℃ in KJ/mol
∆H-∆U=∆ngRT
∆ng=12-15=-3
∆ngRT=(-3)×(8.314×10^-3KJ/mol/K)×298=-7.43